# TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks

### v1

@@ 6 Baseline Agent
7 Experimental Results
In this section, we evaluate popular foundation models, both closed and open, on TheAgentCompany benchmark. We use OpenHands CodeAct agent (§ 6) for all experiments This serves as a baseline for future development of both the foundation LLMs and the agent infrastructure. Note that since LLM evaluators and NPCs are part of the environment rather than the agent being evaulated, we fix their backbone LLM to Claude-3-5-Sonnet-20241022, which demonstrated the best qualitative accuracy in simulating human colleagues and judging deliverables in preliminary experiments.
Model Success Score Steps Costs
API-based Models
Claude-3.5-Sonnet 24.0% 34.4% 29.17 \$6.34
Gemini-2.0-Flash 11.4% 19.0% 39.85 \$0.79
GPT-4o 8.6% 16.7% 14.55 \$1.29
Gemini-1.5-Pro 3.4% 8.0% 22.10 \$6.78
Amazon-Nova-Pro-v1 1.7% 5.7% 19.59 \$1.55
Open-weights Models
Llama-3.1-405b 7.4% 14.1% 22.95 \$3.21
Llama-3.3-70b 6.9% 12.8% 20.93 \$0.93
Qwen-2.5-72b 5.7% 11.8% 23.99 \$1.53
Llama-3.1-70b 1.7% 6.5% 19.18 \$0.83
Qwen-2-72b 1.1% 4.2% 23.70 \$0.28
Model Success (%) Score (%) Success (%) Score (%) Success (%) Score (%) Success (%) Score (%)
API-based Models
Claude-3.5-Sonnet 30.99 40.25 41.18 50.37 21.52 34.68 10.00 21.81
Gemini-2.0-Flash 11.27 18.21 17.65 29.84 13.92 23.34 2.86 8.52
GPT-4o 11.27 19.46 23.53 33.68 5.06 16.08 1.43 7.76
Gemini-1.5-Pro 2.82 3.88 5.88 14.05 3.80 10.97 0.00 4.22
Amazon-Nova-Pro-v1 2.82 7.22 5.88 16.67 1.27 5.36 0.00 2.43
Open-weights Models
Llama-3.1-405b 5.63 11.84 29.41 39.12 8.86 16.46 0.00 4.45
Llama-3.3-70b 8.45 14.26 11.76 21.65 5.06 12.06 0.00 3.76
Qwen-2.5-72b 5.63 11.33 11.76 23.56 5.06 12.60 0.00 4.14
Llama-3.1-70b 1.41 6.09 5.88 15.35 2.53 8.23 0.00 3.32
Qwen-2-72b 1.41 1.94 5.88 12.45 0.00 4.88 0.00 2.60
Model Success Score Success Score Success Score Success Score Success Score Success Score Success Score
API-based Models
Claude-3.5-Sonnet 30.43 38.02 35.71 51.31 14.29 21.70 0.00 11.59 24.14 34.49 8.33 25.17 12.50 22.40
Gemini-2.0-Flash 13.04 18.99 17.86 31.71 0.00 6.49 6.67 15.20 17.24 23.08 0.00 4.31 0.00 10.05
GPT-4o 13.04 19.18 17.86 32.27 0.00 4.70 6.67 13.89 0.00 8.28 0.00 7.36 0.00 10.78
Gemini-1.5-Pro 4.35 5.64 3.57 13.19 0.00 4.82 6.67 9.92 3.45 11.42 0.00 2.78 0.00 8.07
Amazon-Nova-Pro-v1 2.90 6.07 3.57 12.54 0.00 3.27 0.00 0.00 0.00 4.27 0.00 2.78 0.00 2.86
Open-weights Models
Llama-3.1-405b 5.80 11.33 21.43 35.62 0.00 5.42 0.00 3.33 6.90 12.56 0.00 5.00 12.50 17.45
Llama-3.3-70b 11.59 16.49 7.14 19.83 0.00 4.70 0.00 1.67 6.90 11.38 0.00 5.69 0.00 7.03
Qwen-2.5-72b 7.25 11.99 10.71 22.90 0.00 5.42 0.00 2.14 6.90 12.36 0.00 7.15 0.00 5.99
Llama-3.1-70b 1.45 4.77 3.57 15.16 0.00 5.42 0.00 2.42 3.45 7.19 0.00 3.82 0.00 2.86
Qwen-2-72b 2.90 3.68 0.00 7.44 0.00 4.70 0.00 0.56 0.00 4.14 0.00 3.61 0.00 4.95
7.1 Result Overview
Table 3 shows the evaluation results of both closed and open foundation models on the full evaluation set of TheAgentCompany (175 tasks). We can see that the Claude-3.5-Sonnet is the clear winner across all models. However, even with the strongest frontier model, it only manages to complete 24% of the total tasks and achieves a score of 34.4% taking into account partial completion credits. Note that this result comes at a cost: It requires an average of almost 30 steps and more than \$6 to complete each task, making it the most expensive model to run both in time and in cost. This is expected as most of the tasks in our benchmark are of long-horizon nature. The Gemini 2.0 Flash model that comes second in terms of capability requires 40 steps on average to complete the tasks, which is time consuming, yet only to achieve less than half the success rate compared to the top-performing model. Surprisingly, its cost is less than \$1, making it a very cost-efficient yet relatively strong model. A qualitative examination demonstrated that this was due to instances where the agent got stuck in a loop or aimlessly explored the environment.
Among the open-weight models, Llama 3.1 (405B) achieves the highest performance, nearly on par with OpenAI’s GPT-4o model, though still having a big gap behind the leading Claude 3.5 Sonnet. Interestingly, comparing the number of steps and costs between the open Llama 3.1 (405B) model and the closed OpenAI GPT-4o model, Llama 3.1 takes more steps and costs nearly 2x more to run, while having a lower success than GPT-4o. Anecdotally, our inspection showed that GPT-4o seems to be better at giving up early, saving steps and costs if the task is clearly out of the capacity range of the agent. This suggests that open-weight models are not always the most cost-effective choice in agents given serving cost, especially with highly complex tasks.
On the other hand, the newer generation of Llama model, Llama 3.3 (70B) achieves a considerably high performance of 6.9% success rate, on par with the much larger (405B), older generation (Llama 3.1) model. This model also costs significantly less because of its smaller size. This suggests a promising future for LLM development, as smaller and more efficient models begin to catch up in agent performance.
7.2 Analysis
How well do agents operate on different platforms?
Table 4 presents performance breakdown on tasks that involve different platforms in TheAgentCompany. A task is categorized under a platform if one of the platforms that the task requires it. From 4(a), we can see that most models struggle with RocketChat and ownCloud. RocketChat platform is where all the social interaction with peers happens, and the low performance on this platform suggests that current-day LLMs still need improvements in communicating with others. ownCloud platform provides online Office suite functionality, and due to the complexity of the UI of web-based Office software, it is expected that current LLMs fail badly on the platform. This suggests that the browsing capability of the agents, especially on more complex websites, still needs improvement. These results underscore the inherent challenges and complexities of performing tasks that occur in real-world work environments, involve social interaction, or require understanding and navigating complex web interfaces.
(a) Success rate across platforms
(b) Success rate across task categories
Figure 4: Comparing agent success rate across platforms (left) and task categories (right).
How well do agents perform on different type of tasks?
Table 5 presents performance breakdown on different types of tasks in TheAgentCompany. According to the nature of the task, i.e. what kind of professionals are usually assigned with the task, the tasks in TheAgentCompany can be categorized into several job departments: Software Development Engineering (SDE), Project Management (PM), Data Science (DS), Administrative (Admin), Human Resources (HR), Financial (Finance) and all the remaining (Other).
From the success rate demonstrated in 4(b), we can see that data science, administrative, and finance tasks are among the lowest, with many LLMs completing none of the tasks successfully, and even the strongest Claude model achieving much less than the rest of the tasks. On the other hand, software engineering tasks, which may seem like much harder tasks for many humans, result in a higher success rate. This suggests that there exists a gap between the perceived difficulty of the tasks for humans versus the difficulty for LLM agents.
For example, some tasks in the administrative and finance category involves making spreadsheets, collecting and filling in a lot of information from various people, or reading and understanding images scanned by employees. These tasks are arguably easier conceptually for humans in terms of professional skill sets than software engineering, as SDE jobs usually have a higher barrier of entry and more prerequisites for certain knowledge. However, most LLMs achieve a much higher score on the SDE tasks. However, LLMs fail these seemingly easier tasks due to lack of ability to understand documents, communicate with other people, navigate complex software and tedious processes, and autonomously automate repetitive tasks. We hypothesize that part of the reason lies in the fact that current LLM development is heavily based on software engineering abilities, such as coding, due to several high profile benchmarks that measure this capability (e.g. HumanEval, SWE-Bench) as well as the abundance of publicly available training data related to software. On the other hand, administrative and financial tasks, are usually private data within companies, not readily available for training LLMs.
7.3 Common Agent Failures
Lack of commonsense
Some tasks are failed because the agent lacks the common sense and domain background knowledge required to infer implicit assumptions. For example, one task asked the agent to “Write the responses to /workspace/answer.docx” but does not explicitly states that this is a Microsoft Word file. A human can infer this requirement from the file extension. The agent instead treats it as a plain text file, writing text directly to the file, resulting in a task failure.
Lack of social skills
Sometimes, the agent fails to understand the implications and goals in the social conversations with colleagues in TheAgentCompany. For example, one task involves asking Alex for help, and the agent first successfully asks the right question “Could you tell me who I should introduce myself to next on the team?” Then the simulated colleague Alex replied “You should introduce yourself to Chen Xinyi next. She’s on our frontend team and would be a great person to connect with!” At this point, a human would then talk to Chen Xinyi, but instead the agent then decides to not follow up with her, and prematurely considers the task accomplished.
Incompetence in browsing
Often times, the biggest obstacle in tasks is the parts that require browsing the Web. This is expected as browsing is still hard for agents given the complexity of modern-day web UIs and the numerous distractions on a webpage. For example, on many tasks that involve ownCloud, the closable popup that sometimes shows up and asks the user to download the mobile phone apps for a better experience has become an obstacle. Humans can simply click on the ‘x’ to close the popup, while the agents are stuck. Similarly, when trying to download a file from ownCloud, there are several popups to click through before the actual download, and each step is error prone for agents due to the complex UI.
Deceiving oneself
Interestingly, we find that for some tasks, when the agent is not clear what the next steps should be, it sometimes try to be clever and create fake “shortcuts” that omit the hard part of a task. For example, during the execution of one task, the agent cannot find the right person to ask questions on RocketChat. As a result, it then decides to create a shortcut solution by renaming another user to the name of the intended user.
8 Implications and Future Directions
That said, this is just a first step towards forming a firmer grasp on how AI may affect the tasks performed within a workspace, and it has its limitations. First, our tasks are generally on the more straightforward side due to the need to automatically evaluate with programs and test cases, and we do not cover more complex creative tasks such as brainstorming new product ideas or designing system architectures. Second, we are only using one agent scaffold as the baseline performance, and others may differ in performance. Third, while it would be interesting to know the actual performance of human professionals on these tasks to understand how LLM agents perform in comparison, due to resource limitations we were not able to perform this comparison in the current iteration of TheAgentCompany. Fourth, the topic and content of the tasks were mostly created through introspection by people familiar with these workspaces, which may result in some disconnect with actual tasks performed in enterprise settings.
Author Contributions
This work was an open source collaborative effort between multiple institutions and many independent individuals. We used a point-based system to determine contributions and award authorship. Frank Xu, Boxuan Li and Yufan Song led the project, coordinating overall development and paper writing efforts. Detailed contributions were as follows:
  • Task Design: Frank Xu, Yufan Song, Boxuan Li, Zora Wang, Shuyan Zhou, Graham Neubig
  • Infrastructure Development: Yufan Song, Boxuan Li
  • Experiment: Boxuan Li, Frank Xu, Yufan Song
  • Sotopia Integration: Yufan Song, Xuhui Zhou
  • Task Development: Boxuan Li, Yufan Song, Frank Xu, Graham Neubig, Yuxuan Tang, Mengxue Bao, Kritanjali Jain, Zhitong Guo, Murong Cao, Mingyang Yang, Hao Yang Lu, Amaad Martin, Zhe Su, Leander Maben, Raj Mehta, Yiqing Xie, Zora Wang, Xuhui Zhou, Wayne Chi, Lawrence Jang
  • Ideation, Discussion and Formulation: Frank Xu, Shuyan Zhou, Xuhui Zhou, Zora Wang, Wayne Chi, Yufan Song, Boxuan Li, Lawrence Jang, Graham Neubig
  • Advising: Graham Neubig advised the project, providing guidance, resources, and substantial paper editing.
Acknowledgments
This work was supported by a grant from Open Philanthropy. We thank Daniel Fried, Ruslan Salakhutdinov, Chris Donahue, Jing Yu Koh, Brandon Trabucco, Julie Nys, Olga Rogunova, Aditya Soni, Alex Lawsen, and Stephen McAleer for their insightful discussions in various stages of the project. We also thank the OpenHands team for their help on some of the task implementations.
Appendix A More TheAgentCompany Environment Details
A.1 TheAgentCompany Overview
@@ 6. Low-Code Event-Driven Application Platform (based on Node-RED)
A.2 TheAgentCompany Employee Roster with Project Assignments and Slack Channels
@@ 18. Jessica Chen (Female, 28 years old)
A.3 TheAgentCompany Q3 2024 Quarterly Sprint Goals
@@ 6. Low-Code Platform Team (Node-RED)
A.4 TheAgentCompany Internal Documents
@@
Appendix B Agent-Simulated Colleagues Conversation Examples

### v2

@@ Appendix D Baseline Agent Details
Appendix E LLM-as-a-Judge
In TheAgentCompany, there are 51 tasks (29%) involving LLM evaluation. LLM-based evaluators are mainly used in well-defined tasks that require simple information extraction and classification, which has been shown to have high precision  Zheng et al. (2023). Furthermore, we first use deterministic keyword matches and then use LLM as a fallback. Our use of LLM evaluation is a supplement to the deterministic evaluator, rather than a replacement. In addition, the assessors were reviewed and tested by 3 to 5 contributors and continuous integration was carried out to ensure robustness. TheAgentCompany scoring framework includes both step-by-step and final-result evaluations. In cases where a correct and sound deliverable receives no credit due to an evaluator misjudging an intermediate step, the scoring framework would override the score and grant full credit. Above all, the main concerns about the introduction of LLMs as judges should be dismissed.
Appendix F LLM-as-Colleagues
Although we acknowledge that LLMs can make mistakes when acting in different roles, we empirically found such errors to be rare. A careful examination of the early version of all 41 tasks that simulated colleagues involved revealed only one trajectory with an issue that stemmed from role-play prompt ambiguity, which was already fixed in the latest TheAgentCompany version. The prompt to define each role (i.e. the NPC) is simple and straightforward without prompt engineering. As more powerful models are released to act as simulated colleagues, the likelihood of errors decreases.
For NPCs, most context interactions occur within a few hundred tokens. If we assume 3000 input tokens and 1000 output tokens, which is a reasonable upper bound, then according to Claude 3.5 Sonnet’s pricing (Input token price: \$3.00, Output token price: \$15.00 per 1M tokens, assuming no prompt caching), the cost per NPC interaction would not exceed \$0.024, and in most cases, it is much lower.
Appendix G More TheAgentCompany Environment Details
G.1 TheAgentCompany Overview
@@ 6. Low-Code Event-Driven Application Platform (based on Node-RED)
G.2 TheAgentCompany Employee Roster with Project Assignments and Slack Channels
@@ 18. Jessica Chen (Female, 28 years old)
G.3 TheAgentCompany Q3 2024 Quarterly Sprint Goals
@@ 6. Low-Code Platform Team (Node-RED)
G.4 TheAgentCompany Internal Documents
@@
Appendix H Agent-Simulated Colleagues Conversation Examples
@@
Appendix I Detailed Experimental Results
Table 4 presents performance breakdown on tasks that involve different platforms in TheAgentCompany. A task is categorized under a platform if it is one of the platforms for which the task requires it. Table 5 presents the performance breakdown for different types of tasks in TheAgentCompany. Depending on the nature of the task, i.e. what kind of professionals are usually assigned to the task, the tasks in TheAgentCompany can be categorized into various departments of jobs. Software Development Engineering (SDE), Project Management (PM), Data Science (DS), Administrative (Admin), Human Resources (HR), Financial (Finance) and all the remaining (Other).
Agent Model Success (%) Score (%) Success (%) Score (%) Success (%) Score (%) Success (%) Score (%)
API-based Models
OpenHands 0.28.1 Gemini-2.5-Pro 33.80 43.36 41.18 51.67 29.11 40.39 12.86 22.44
OpenHands 0.28.1 Claude-3.7-Sonnet 23.94 32.60 41.18 52.63 29.11 41.81 11.43 23.97
OpenHands 0.14.2 Claude-3.5-Sonnet 30.99 40.25 41.18 50.37 21.52 34.68 10.00 21.81
OpenHands 0.14.2 Gemini-2.0-Flash 11.27 18.21 17.65 29.84 13.92 23.34 2.86 8.52
OpenHands 0.14.2 GPT-4o 11.27 19.46 23.53 33.68 5.06 16.08 1.43 7.76
OWL Roleplay GPT-4o & o3-mini 5.63 12.51 5.88 15.39 3.80 11.07 0.00 5.85
OpenHands 0.14.2 Gemini-1.5-Pro 2.82 3.88 5.88 14.05 3.80 10.97 0.00 4.22
OpenHands 0.14.2 Amazon-Nova-Pro-v1 2.82 7.22 5.88 16.67 1.27 5.36 0.00 2.43
Open-weights Models
OpenHands 0.14.2 Llama-3.1-405b 5.63 11.84 29.41 39.12 8.86 16.46 0.00 4.45
OpenHands 0.14.2 Llama-3.3-70b 8.45 14.26 11.76 21.65 5.06 12.06 0.00 3.76
OpenHands 0.14.2 Qwen-2.5-72b 5.63 11.33 11.76 23.56 5.06 12.60 0.00 4.14
OpenHands 0.14.2 Llama-3.1-70b 1.41 6.09 5.88 15.35 2.53 8.23 0.00 3.32
OpenHands 0.14.2 Qwen-2-72b 1.41 1.94 5.88 12.45 0.00 4.88 0.00 2.60
Agent Model Success Score Success Score Success Score Success Score Success Score Success Score Success Score
API-based Models
OpenHands 0.28.1 Gemini-2.5-Pro 37.68 45.05 39.29 52.61 14.29 20.06 13.33 19.17 34.48 44.98 8.33 21.56 12.50 20.16
OpenHands 0.28.1 Claude-3.7-Sonnet 30.43 36.78 42.86 56.32 14.29 22.80 13.33 23.89 27.59 40.02 0.00 18.89 12.50 24.06
OpenHands 0.14.2 Claude-3.5-Sonnet 30.43 38.02 35.71 51.31 14.29 21.70 0.00 11.59 24.14 34.49 8.33 25.17 12.50 22.40
OpenHands 0.14.2 Gemini-2.0-Flash 13.04 18.99 17.86 31.71 0.00 6.49 6.67 15.20 17.24 23.08 0.00 4.31 0.00 10.05
OpenHands 0.14.2 GPT-4o 13.04 19.18 17.86 32.27 0.00 4.70 6.67 13.89 0.00 8.28 0.00 7.36 0.00 10.78
OWL RolePlay GPT-4o & o3-mini 5.80 11.67 3.57 16.19 0.00 4.82 0.00 3.53 6.90 14.30 0.00 9.58 0.00 2.86
OpenHands 0.14.2 Gemini-1.5-Pro 4.35 5.64 3.57 13.19 0.00 4.82 6.67 9.92 3.45 11.42 0.00 2.78 0.00 8.07
OpenHands 0.14.2 Amazon-Nova-Pro-v1 2.90 6.07 3.57 12.54 0.00 3.27 0.00 0.00 0.00 4.27 0.00 2.78 0.00 2.86
Open-weights Models
OpenHands 0.14.2 Llama-3.1-405b 5.80 11.33 21.43 35.62 0.00 5.42 0.00 3.33 6.90 12.56 0.00 5.00 12.50 17.45
OpenHands 0.14.2 Llama-3.3-70b 11.59 16.49 7.14 19.83 0.00 4.70 0.00 1.67 6.90 11.38 0.00 5.69 0.00 7.03
OpenHands 0.14.2 Qwen-2.5-72b 7.25 11.99 10.71 22.90 0.00 5.42 0.00 2.14 6.90 12.36 0.00 7.15 0.00 5.99
OpenHands 0.14.2 Llama-3.1-70b 1.45 4.77 3.57 15.16 0.00 5.42 0.00 2.42 3.45 7.19 0.00 3.82 0.00 2.86
OpenHands 0.14.2 Qwen-2-72b 2.90 3.68 0.00 7.44 0.00 4.70 0.00 0.56 0.00 4.14 0.00 3.61 0.00 4.95
Appendix J Limitation
During the construction of this benchmark, we had people test and complete all tasks to ensure the feasibility of the tasks and the evaluator. However, we did not collect human performance data. The main reason is cost, as completing a single task can take anywhere from 10 minutes to several hours, and currently we do not have the means to recruit a large number of people with the relevant background to complete the tasks on a scale. However, this does not affect the evaluation of the current models in the benchmark.